# Hardware and Software Platform Inference

### v1

@@ 6.5 Robustness of HSPI
For logits, introducing random bit flips into the decision-making process can help mask the unique quantization patterns associated with specific hardware architectures. This technique adds a layer of noise that obscures the underlying hardware fingerprints, while incurring minimal overheads associated with sampling random numbers. Similarly, for border inputs, adding random noise to the input data can disrupt the precise calculations that lead to divergent behavior across different hardware and software configurations. This noise injection makes it harder for attackers to identify the specific conditions that trigger variations in model outputs, yet can potentially come with utility degradation.
7 Conclusion
In this paper, we introduce Hardware and Software Platform Inference (HSPI) – a method for identifying the underlying hardware and software stack solely based on the input-output behavior of machine learning models. Our approach leverages inherent quantization limitations and the arithmetic operation orders of different GPU architectures to distinguish between various GPU types and software stacks. By analyzing quantization patterns in the model’s outputs, our classification framework can accurately identify the specific GPU and software configuration used for model inference. Our findings demonstrate the feasibility of inferring this information from black-box models, underscoring its implications for ensuring transparency and accountability in the LLM market.
Appendix A Appendix
A.1 Fine-tuning Hyper-parameters of Vision Models
The checkpoints of vision models in Section 5 are downloaded from torchvision 333 https://pytorch.org/vision/stable/models.html#classification. We fine-tune these models using the following hyper-parameters:
Optimizer Learning Rate LR Scheduler Batch size Num Epochs SGD 1e-3 Linear 128 3
A.2 Quantization and Test Accuracy of Vision Models
We perform a sanity check before the HSPI experiments in case the quantization breaks the models.

### v2

@@ Appendix F Robustness of HSPI
Several mitigation strategies can be employed to make the inference of hardware and software platforms significantly more difficult. These strategies focus on disrupting the patterns and subtle variations exploited by HSPI. For logits, introducing random bit flips into the decision-making process can help mask the unique quantization patterns associated with specific hardware architectures. This technique adds a layer of noise that obscures the underlying hardware fingerprints, while incurring minimal overheads associated with sampling random numbers. Similarly, for border inputs, adding random noise to the input data can disrupt the precise calculations that lead to divergent behavior across different hardware and software configurations. This noise injection makes it harder for attackers to identify the specific conditions that trigger variations in model outputs, yet can potentially come with utility degradation.